Y=xdy/dx+1/2(dy/dx)^2
1
Expert's answer
2020-08-16T20:32:13-0400

As per the question,

y=x"\frac{dy}{dx}"+"\frac{d^2y}{2dx^2}"

Arranging the above equation as,

"\frac{d^2y}{dx^2}" +2x"\frac{dy}{dx}" -2y=0 .....(1)(say)

let y=xt,

we take y=xt since It will fulfill the condition

since we have to reduce Second order differential equation into first order.

differentiate both sides with respect to x,

"\frac{dy}{dx}" =t

Again differentiating above equation with respect to x,

"\frac{d^2y}{dx^2}" ="\frac{dt}{dx}"

Putting the above values in equation 1 we get,

"\frac{dt}{dx}" +2xt-2xt=0


"\frac{dt}{dx}" =0

so t=constant=k(say)

since t="\frac{y}{x}"

"so \frac{y}{x}" =k

y=kx

This is the required solution.